OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Beluga Fun Tours: Age Ticket Price Under 3 years free 3 to 12 years $34 Over 12 years $45 On a certain day, the company took 103 people on whale watching trips. There were 64 children aged 12 and under, of which some children were under 3 years. If x represents the number of children under 3 years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (anonymous):

(64 - x)34 + (103 - 64) ⋅ 45 = C (103 + x)34 + (103 -(64 + x)) ⋅ 45 = C (103 - x)34 + (103 - 64) ⋅ 45 = C (64 + x)34 + (103 -(64 + x)) ⋅ 45 = C

OpenStudy (anonymous):

I highly believe its a but just checking

OpenStudy (doc.brown):

Don't believe. Know.

OpenStudy (doc.brown):

Yeah, you're right.

OpenStudy (mathmale):

SectorZ423: How about explaining your reasoning and showing some of the work you've done?

OpenStudy (anonymous):

Well there is a possible 64 children that were under 4 (3 and under) so I just went with my gut and took out B and C and worked on A and D and since D is adding more of the children that were already there then I just kinda took A as the answer

OpenStudy (mathmale):

I'd suggest you make a table of the 3 groups of people and assign a letter to each to represent the number of people in each group: kids < 3 x kids between 3 and 12 y people over 12 z We already know that x represents the number of kids under 3; this is given. We need to express y and z in terms of x. Note that the total number of kids 12 and under is 64, so then the number who are between 3 and 12 is y = 64 - x. Can you come up with an expression in x alone for the number of people over 12? I'm not saying a "gut feeling" won't work or isn't good, but a systematic approach improves your chances of obtaining the correct answer.

OpenStudy (mathmale):

kids < 3 x Leave as is kids between 3 and 12 y = 64-x Replace y by 64-x people over 12 z = ????

OpenStudy (anonymous):

Hmm so I was right?

OpenStudy (anonymous):

If I was sorry im not all that good at explaining things sometimes in my mind I figure it out without noticing and thats how I get the "gut feeling"